# GARLIC: GAussian Representation LearnIng for spaCe partitioning

### v1

@@ 2.2 Optimization Objective
We adapt ideas from 3D Gaussian Splatting (3DGS) [29] to high-dimensional nearest neighbor search. While 3DGS performs well in 3D scene modeling, its reliance on Structure-from-Motion (SfM) initialization and view-space heuristics breaks down in high dimensions, where geometry is abstract and dimensions encode heterogeneous features. Splitting and cloning based on spatial cues become unreliable, and optimization suffers in sparse regions. These issues are exacerbated by the curse of dimensionality and clustering, which degrade distance metrics and make balanced partitioning ineffective. High-dimensional indexing demands adaptive, data-aware refinement, something 3DGS heuristics cannot provide.
@@
To handle these issues, inspired by Self-Supervised learning techniques [9, 45], that use information theoretic objectives to increase the capacity of the representation space, we introduce similar information objectives to guide optimization towards a more informative and balanced solution without excessive prior usage. More specifically, we introduce a divergence-based objective that acts as a reconstruction loss, and regularize it to prevent information explosion (i.e., uncontrolled growth and excessive overlap of Gaussians). The divergence loss $\mathcal{L}_{\text{div}}$ is defined as:
$\mathcal{L}_{\text{div}}=\frac{1}{N}\sum_{\mathbf{x}\in X}\left(\min_{g_{i}\in\mathcal{G}}\delta_{M}(\mathbf{x},g_{i})-\tau\right)^{+},$
where $\delta_{M}(\mathbf{x},g_{i})=||\mathbf{L}_{i}^{-1}(\mathbf{x}-\boldsymbol{\mu}_{i})||_{2}$ denotes the Mahalanobis distance, $(\cdot)^{+}=\max(0,\cdot)$ and $\tau$ is a standard deviation threshold. It penalizes points that fall outside a Gaussian’s coverage radius ($\delta_{M}(\mathbf{x}_{i},g_{i})>\tau$), encouraging Gaussians to expand and cover these points, while points inside ($\delta_{M}(\mathbf{x}_{i},g_{i})\leq\tau$) do not contribute, allowing controlled expansion but preventing infinite growth.
@@
Still, Gaussians can overlap, leading to redundant information. To surpass this issue, and mitigate fuzzy assignments of points to Gaussians, we introduce a covariance-based regularization, which encourages each Gaussian to dominate its assigned points. Specifically, given a point $\mathbf{x}$, we define its coverage set, i.e., the set of Gaussians that satisfy the coverage radius constraint, as $\mathcal{M}(\mathbf{x})=\{g_{i}\in\mathcal{G}\mid\delta_{M}(\mathbf{x},g_{i})\leq\tau\}.$
Then, we compute the normalized soft-assignment probabilities based on Euclidean distances ($\delta_{\mathcal{E}}$) as $p_{i}(\mathbf{x})=e^{-\delta_{\mathcal{E}}(\mathbf{x},\boldsymbol{\mu}_{i})}/\sum_{g_{j}\in\mathcal{M}(\mathbf{x})}e^{-\delta_{\mathcal{E}}(\mathbf{x},\boldsymbol{\mu}_{j})}+\epsilon,\ \forall g_{i}\in\mathcal{M}(\mathbf{x})$. The covariance loss $\mathcal{L}_{\text{cov}}$ is defined as:
$\mathcal{L}_{\text{cov}}=1-\frac{1}{N}\sum_{\mathbf{x}\in X}\max_{g_{i}\in\mathcal{M}(\mathbf{x})}p_{i}(\mathbf{x})$
@@
This loss encourages highly confident (low-entropy) assignments, thereby reducing ambiguity and stabilizing optimization. To prevent excessive expansion of Gaussians caused by the divergence objective and ensure that each Gaussian accurately aligns with its assigned points, we introduce the anchor loss $\mathcal{L}_{\text{anchor}}$:
$\mathcal{L}_{\text{anchor}}=\frac{1}{d|G|}\sum_{g_{i}\in G}\left(||\boldsymbol{\mu}_{i}-\hat{\boldsymbol{\mu}}_{i}||_{2}^{2}+\alpha||\mathbf{L}_{i}{\mathbf{L}_{i}}^{\top}-\hat{\boldsymbol{\Sigma}}_{i}||_{F}^{2}\right),$
@@
where $\hat{\boldsymbol{\mu}}_{i}$ and $\hat{\boldsymbol{\Sigma}}_{i}=\mathrm{Cov}(x\in B_{g_{i}})$ are the empirical mean and covariance of points assigned to Gaussian $g_{i}$, and $\alpha$ is a hyperparameter balancing position and shape. This loss constrains Gaussians by anchoring them closely to their local data distributions, counteracting uncontrolled growth from other loss terms and maintaining geometric alignment and interpretability. Finally, our loss function is defined as:
@@
(b) Grad flow: ${\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathcal{L}_{\text{div}}}$, ${\color[rgb]{0.72,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.72,0,0}\pgfsys@color@cmyk@stroke{0}{0.89}{0.94}{0.28}\pgfsys@color@cmyk@fill{0}{0.89}{0.94}{0.28}\mathcal{L}_{\text{cov}}}$, ${\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\mathcal{L}_{\text{anchor}}}$
Figure 2: Gaussian refinement and associated loss gradients.
@@ 2.3 Adaptive Refinement
Controlling the number of Gaussians during training is key for adapting to data complexity; allocating more in dense regions and fewer in simpler ones. We employ an adaptive refinement strategy that adjusts the Gaussian set based on local point density. Prior approaches rely on positional gradients to guide refinement [29], but these become sparse and unreliable in high dimensions. In our case, we also require search efficiency, so we refine Gaussians whose bucket cardinality exceeds a threshold: $|\mathcal{B}_{g}|>\gamma\cdot|X|$. Splitting is performed by applying DBSCAN [17] to detect two dense clusters within the bucket. If unsuccessful, we fall back to K-Means to produce new means $\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2}$, while covariances are scaled down as $\mathbf{L}_{1}=\mathbf{L}_{2}=\alpha\cdot\mathbf{L}$, with $\alpha<1$.
For cloning, we target Gaussians with a high ratio of outside points to interior ones $\mathcal{B}_{g_{i}}^{in}=\{\mathbf{x}\in X:\delta_{M}(\mathbf{x},g_{i})\leq\tau\}$. Instead of using all nearest outside points, we focus on a boundary region, and select $k=\rho\cdot|\mathcal{B}_{g}^{out}|$ points, where $\rho\in(0,1)$ is a sampling ratio and $\mathcal{B}_{g_{i}}^{out}=\{\mathbf{x}\in X:\tau<\delta_{M}(\mathbf{x},g_{i})<e\tau\text{ and }g_{i}=\arg\min_{g_{j}\in G}\delta_{M}(\mathbf{x},g_{j})\}$ with $e>1$ is the set of all boundary points. The sampled subset $S=\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{k}\}\subset\mathcal{B}_{g}^{out}$ is chosen randomly without replacement. This sampling approach serves several purposes, reducing computational cost, as density estimation in high dimensions is expensive, providing statistical robustness by focusing on representative points and helping to avoid outliers that might exist in the boundary region as Fig. 2 showcases.
The candidate Gaussians are picked with regards to the ration of boundary to interior points: $|\mathcal{B}_{g}^{out}|/|\mathcal{B}_{g}^{in}|>\beta$. The center of the new Gaussian is placed at the point with the highest local density in the boundary region, such that for a point $\mathbf{p}$ in the boundary region, we compute its local density as:
$\rho(\mathbf{p})=\frac{1}{\frac{1}{k}\sum_{j=1}^{k}\delta_{\mathcal{E}}(\mathbf{p},\text{NN}_{j}(\mathbf{p}))},$
here $\text{NN}_{j}(\mathbf{p})$ is the $j$-th nearest neighbor of $\mathbf{p}$ and $\delta_{\mathcal{E}}(\cdot,\cdot)$ is the Euclidean distance. We select the point with the highest density: $\mathbf{p}^{*}=\arg\max_{\mathbf{p}}\rho(\mathbf{p})$. The covariance matrix is then cloned, with $\mathbf{L}_{new}=\mathbf{L}$.
@@ 2.4 Quantization
After optimizing our Gaussians, we assign points that fall outside the coverage radius to the nearest Gaussian such that our final buckets are:
$B_{g_{i}}=\{\mathbf{x}\in X\mid\delta_{M}(\mathbf{x},g_{i})\leq\tau\}\cup\{\mathbf{x}\in X\mid\delta_{M}(\mathbf{x},g_{i})>\tau\text{ and }g_{i}=\arg\min_{g_{j}\in G}\delta_{M}(\mathbf{x},g_{j})\}$
Algorithm 1 Build Index
1:Input: Dataset $\mathcal{D}$ of $d$-dimensional points
2:Initialize Gaussians on $\mathcal{D}$
3:$\mathcal{G}\leftarrow$ Obtain $K$ Gaussians $g_{i}=\mathcal{N}(\mu_{i},L_{i})$ (Sec. 2.22.3)
4:for each point $p$ do
5:    Assign $p$ to all Gaussians where $d_{M}(p,g_{i})<\tau$, $\forall i\in[K]$
6:    if $p$ not assigned to any Gaussian then
7:         Assign $p$ to nearest Gaussian w.r.t. $d_{M}$
8:    end if
9:end for
10:for each Gaussian $g_{i}\in\mathcal{G}$ do
11:    $\mathcal{Q}_{i}\leftarrow$ Quantize indexed data of $g_{i}$ (Sec. 2.4)
12:end for
13:return $\{(g_{i},\mathcal{Q}_{i})\}_{i\in[K]}$
While our Gaussian-based buckets provide an effective organization of the data space, exhaustive search within buckets becomes prohibitively expensive as dimensionality increases. For high-dimensional nearest neighbor search, we must avoid calculating distances to all points within candidate buckets. Standard approximate search methods, however, often do not account for the unique structure provided by our Gaussian-based partitioning.
We introduce a quantization scheme that leverages the inherent properties of our Gaussian model. The key insight is that once we have assigned points to appropriate Gaussian buckets, we can apply efficient Euclidean-based quantization within each bucket’s local coordinate system, reducing the number of distance computations required at query time.
Algorithm 2 Query Index
1:Input: Index structure set $\{(g_{i},\mathcal{Q}_{i})\}_{i\in[K]}$, query $q$
2:$g_{i}\leftarrow\arg\min_{i:g_{i}\in\mathcal{G}}d_{M}(q,g_{i})$
3:$B\leftarrow$ Traverse subset of bins from $\mathcal{Q}_{i}$ (Sec. 2.4)
4:Measure $L_{2}$ distance to candidates in $B$
5:Terminate if condition (Sec. 2.5)
6:return nearest neighbor of $q$ in $B$
@@
For each bucket $B_{g}$, we apply PCA to reduce dimensionality while preserving the local structure $\mathbf{P}_{g}=\text{PCA}(\{\mathbf{x}-\bar{\mathbf{x}}_{g}\mid\mathbf{x}\in B_{g}\},r)$ where $\bar{\mathbf{x}}_{g}$ is the mean of points in bucket $B_{g}$, $r$ is the reduced dimensionality (typically $r\ll d$, constant in practice), and $\mathbf{P}_{g}\in\mathbb{R}^{d\times r}$ contains the principal components. Each point $\mathbf{x}\in B_{g}$ is projected into this lower-dimensional space $\mathbf{x}^{r}=\mathbf{P}_{g}^{T}(\mathbf{x}-\bar{\mathbf{x}}_{g})$, after which its transformed to hyperspherical coordinates $\mathbf{s}=\text{cart2sph}(\mathbf{x}^{r})$, where the transformation for an $r$-dimensional point yields coordinates $\mathbf{s}=(s_{1},s_{2},\ldots,s_{r})$.
The hyperspherical space is then partitioned into bins defined by $n_{r}$ radial divisions and $n_{a}$ angular divisions per dimension: $\mathcal{B}_{g}=\{B_{i,\mathbf{j}}\mid i\in\{1,\ldots,n_{r}\},\mathbf{j}\in\{1,\ldots,n_{a}\}^{r-1}\}$ with boundaries $r_{i}=r_{\min}+(r_{\max}-r_{\min})\cdot i/n_{r},\ i\in\{0,\ldots,n_{r}\}$ and $\theta_{j,k}=\theta_{\min,k}+(\theta_{\max,k}-\theta_{\min,k})\cdot j/n_{a},\ j\in\{0,\ldots,n_{a}\},k\in\{1,\ldots,r-1\}$. Alg. 1 illustrates the construction process of the index structure.
@@ 2.5 Inference
At query time, for a query point $\mathbf{q}$, we employ a multi-stage approach to efficiently retrieve approximate nearest neighbors. First, we identify the $k$ nearest Gaussians using the Mahalanobis distance. For each selected Gaussian bucket $g$, we project the query into the bucket’s local PCA space, $\mathbf{q}^{r}_{g}=\mathbf{P}_{g}^{T}(\mathbf{q}-\bar{\mathbf{x}}_{g})$, and convert it to hyperspherical coordinates $\mathbf{s}_{g}=\text{cart2sph}(\mathbf{q}^{r}_{g})$. We then compute the distances to all bins in the bucket and sort them in ascending order. For each bin $B_{i,\mathbf{j}}$ with boundaries defined by $r$-intervals $[r_{i},r_{i+1}]$ and angular intervals $[\theta_{j,k},\theta_{j+1,k}]$, we compute the distance $d(\mathbf{s}_{g},B_{i,\mathbf{j}})=\min_{\mathbf{y}\in B_{i,\mathbf{j}}}\|\mathbf{s}_{g}-\mathbf{y}\|_{2}.$
This minimization problem is efficiently solved by first checking if $\mathbf{s}_{g}$ is inside the bin. If it is outside, we solve the constrained optimization problem by initializing a point $\mathbf{y}^{0}$ with its radial component clipped to $[r_{\min},r_{\max}]$ and angular components set to the midpoints of their respective intervals. We then minimize $\|\mathbf{s}_{g}-\mathbf{y}\|_{2}$ subject to the bin boundary constraints using the L-BFGS-B [10] algorithm, which respects the bounds $r_{\min}\leq y_{1}\leq r_{\max}$ and $\theta_{\min,j}\leq y_{j+1}\leq\theta_{\max,j}$ for the radial and angular components respectively.
We then search for points in bins sequentially according to their sorted distances, computing exact Euclidean distances between the query and the points in each bin. The search terminates under the condition that we have examined a predefined percentage of bins (typically 30%). Given a query, Alg. 2 showcases the procedure of retrieving the nearest neighbor using our index structure.
3 Experiments
Datasets. We evaluate our method on three benchmark datasets: SIFT1M [37] (128-dimensional image descriptors with one million points), MNIST [32] and Fashion-MNIST [44] (784-dimensional vectors from 28×28 grayscale images). These datasets provide diverse characteristics in terms of extrinsic and intrinsic dimensionality, scale, and data distribution, offering a comprehensive evaluation framework in Euclidean space. Additional information about the datasets and their licenses are provided in the Supplementary Material (SM).
@@ 3 Experiments
Ours 10%Ours 5%Ours 2.5%Ours 1%
(a) Efficiency
(b) MNIST
(c) Fashion-MNIST
@@
(e) SIFT1M
(f) MNIST
(g) Fashion-MNIST

### v2

@@ 2.2 Optimization Objective
High-dimensional indexing requires objectives that remain stable under the curse of dimensionality and robust to heterogeneous feature distributions. Classical partitioning heuristics (e.g., balanced splits, uniform clustering) degrade in such regimes: dense regions are fragmented, sparse regions are merged, and distance metrics lose discriminative power Aggarwal et al. (2001), Aumüller et al. (2020). To overcome this, we draw inspiration from self-supervised learning methods such as VICReg Bardes et al. (2022) and Barlow Twins Zbontar et al. (2021), which employ information-theoretic objectives to increase representational capacity without supervision. Analogously, we introduce objectives that guide Gaussians to (i) cover the data distribution, (ii) assign points with high confidence, and (iii) remain anchored to local structure. This replaces heuristic spatial rules with principled, differentiable criteria suited to high-dimensional retrieval.
@@
More specifically, we introduce a divergence-based objective that acts as a reconstruction loss, and regularize it to prevent information explosion (i.e., uncontrolled growth and excessive overlap of Gaussians). The divergence loss $\mathcal{L}_{\text{div}}$ is defined as:
$\mathcal{L}_{\text{div}}=\frac{1}{N}\sum_{\mathbf{x}\in X}\left(\min_{g_{i}\in\mathcal{G}}\delta_{M}(\mathbf{x},g_{i})-\tau\right)^{+},$ (1)
[…]
(b) Grad flow: $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{L}_{\text{div}}$}, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{L}_{\text{cov}}$}, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{L}_{\text{anchor}}$}
Figure 2: Gaussian refinement and associated loss gradients.
@@
Still, Gaussians can overlap, leading to redundant information and performance loss. To surpass this issue, and mitigate fuzzy assignments of points to Gaussians, we introduce a covariance-based regularization, which encourages each Gaussian to dominate its assigned points. Specifically, given a point $\mathbf{x}$, we define its coverage set, i.e., the set of Gaussians that satisfy the coverage radius constraint, as $\mathcal{M}(\mathbf{x})=\{g_{i}\in\mathcal{G}\mid\delta_{M}(\mathbf{x},g_{i})\leq\tau\}.$ Then, we compute the normalized soft-assignment probabilities based on Euclidean distances ($\delta_{\mathcal{E}}$) as $p_{i}(\mathbf{x})=e^{-\delta_{\mathcal{E}}(\mathbf{x},\boldsymbol{\mu}_{i})}/\sum_{g_{j}\in\mathcal{M}(\mathbf{x})}e^{-\delta_{\mathcal{E}}(\mathbf{x},\boldsymbol{\mu}_{j})}+\epsilon,\ \forall g_{i}\in\mathcal{M}(\mathbf{x})$. The covariance loss $\mathcal{L}_{\text{cov}}$ is defined as:
$\mathcal{L}_{\text{cov}}=1-\frac{1}{N}\sum_{\mathbf{x}\in X}\max_{g_{i}\in\mathcal{M}(\mathbf{x})}p_{i}(\mathbf{x})$ (2)
This loss encourages highly confident (low-entropy) assignments, thereby reducing ambiguity and stabilizing optimization.
@@
To further prevent excessive expansion of Gaussians caused by the divergence objective and ensure that each Gaussian aligns with its assigned points, we introduce the anchor loss $\mathcal{L}_{\text{anchor}}$:
[…]
@@
where $\hat{\boldsymbol{\mu}}_{i}$ and $\hat{\boldsymbol{\Sigma}}_{i}=\mathrm{Cov}(x\in\mathcal{B}_{g_{i}})$ are the empirical mean and covariance of points assigned to Gaussian $g_{i}$, and $\alpha$ is a hyperparameter balancing position and shape. This loss constrains Gaussians by anchoring them closely to their local data distributions, restraining uncontrolled growth from other loss terms and maintaining geometric alignment. Finally, our loss function is defined as:
@@ 2.3 Adaptive Refinement
Controlling the number of Gaussians during training is key for adapting to data complexity; allocating more in dense regions and fewer in simpler ones. We employ an adaptive refinement strategy that adjusts the Gaussian set based on local point density. Prior approaches rely on positional gradients to guide refinement (Kerbl et al., 2023), but these become sparse and unreliable in high dimensions. In our case, we also require search efficiency, so we refine Gaussians whose cell cardinality exceeds a threshold: $|\mathcal{B}_{g}|>\gamma\cdot|X|$. Splitting is performed by applying clustering (Arthur & Vassilvitskii, 2007) to produce new means $\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2}$, while covariances are scaled down as $\mathbf{L}_{1}=\mathbf{L}_{2}=\alpha\cdot\mathbf{L}$, with $\alpha<1$.
For cloning, we target Gaussians with a high ratio of outside points to interior ones $\mathcal{B}_{g_{i}}^{in}=\{\mathbf{x}\in X:\delta_{M}(\mathbf{x},g_{i})\leq\tau\}$. Instead of using all nearest outside points, we focus on a boundary region, and select $k=\rho\cdot|\mathcal{B}_{g}^{out}|$ points, where $\rho\in(0,1)$ is a sampling ratio and $\mathcal{B}_{g_{i}}^{out}=\{\mathbf{x}\in X:\tau<\delta_{M}(\mathbf{x},g_{i})<e\tau\text{ and }g_{i}=\arg\min_{g_{j}\in\mathcal{G}}\delta_{M}(\mathbf{x},g_{j})\}$ with $e>1$ is the set of all boundary points. The sampled subset $S=\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{k}\}\subset\mathcal{B}_{g}^{out}$ is chosen randomly without replacement. This sampling approach serves several purposes, reducing computational cost, as density estimation in high dimensions is expensive, providing statistical robustness by focusing on representative points and helping to avoid outliers that might exist in the boundary region as Figure 2 showcases. The candidate Gaussians are picked with regard to the ratio of boundary to interior points: $|\mathcal{B}_{g}^{out}|/|\mathcal{B}_{g}^{in}|>\beta$. The center of the new Gaussian is placed at the point with the highest local density in the boundary region, such that for a point $\mathbf{p}$ in the boundary region, we compute its local density as the inverse mean Euclidean distance from the $3$-NN and select the point with the highest density: $\mathbf{p}^{*}=\arg\max_{\mathbf{p}}\rho(\mathbf{p})$. The covariance matrix is then cloned, with $\mathbf{L}_{new}=\mathbf{L}$. In our method, splitting is prioritized over cloning: when a Gaussian grows too large, we first partition it to reduce cell cardinality and maintain search efficiency. Cloning is applied only if splitting is not triggered, serving to refine coverage near dense boundary regions without inflating candidate counts unnecessarily.
@@ 2.4 Quantization
After optimizing our Gaussians, we assign points that fall outside the coverage radius to the nearest Gaussian such that our final cells are: $\mathcal{B}_{g_{i}}=\{\mathbf{x}\in X\mid\delta_{M}(\mathbf{x},g_{i})\leq\tau\}\cup\{\mathbf{x}\in X\mid\delta_{M}(\mathbf{x},g_{i})>\tau\text{ and }g_{i}=\arg\min_{g_{j}\in G}\delta_{M}(\mathbf{x},g_{j})\}$. While Gaussian cells provide an effective partition of the space, brute-force search within them is still prohibitive in high dimensions. Standard ANN methods usually quantize globally, ignoring the anisotropy and local geometry already captured by our Gaussians. We instead introduce a localized quantization scheme: once points are assigned to cells, each cell is treated in its own coordinate system, where Euclidean structure is better aligned with the underlying data. Quantizing in this local basis reduces the number of distance computations per query, while preserving the geometry captured by Mahalanobis distance.
@@
For each cell $\mathcal{B}_{g}$, we apply PCA to reduce dimensionality while preserving the local structure $\mathbf{P}_{g}=\text{PCA}(\{\mathbf{x}-\bar{\mathbf{x}}_{g}\mid\mathbf{x}\in\mathcal{B}_{g}\},r)$ where $\bar{\mathbf{x}}_{g}$ is the mean of points in cell $\mathcal{B}_{g}$, $r$ is the reduced dimensionality (typically $r\ll d$, constant in practice), and $\mathbf{P}_{g}\in\mathbb{R}^{d\times r}$ contains the top-$r$ principal components. This step both lowers computational cost and aligns the local coordinate system with the main variance directions of the data. Each point $\mathbf{x}\in\mathcal{B}_{g}$ is then projected into this subspace $\mathbf{x}^{r}=\mathbf{P}_{g}^{\top}(\mathbf{x}-\bar{\mathbf{x}}_{g})$, which embeds the data in $\mathbb{R}^{r}$ while preserving Euclidean structure up to the discarded components. Finally, we convert $\mathbf{x}^{r}$ into hyperspherical coordinates $\mathbf{s}=\text{cart2sph}(\mathbf{x}^{r})=(s_{1},s_{2},\ldots,s_{r})$, where $s_{1}$ denotes the radial component $\|\mathbf{x}^{r}\|_{2}$ and $s_{2},\dots,s_{r}$ are angular coordinates. This reparameterization retains the Euclidean metric but enables partitioning along radial and angular directions.
The hyperspherical space is partitioned into bins $\mathcal{Q}_{g}=\{B_{i,\mathbf{j}}\mid i\in\{1,\ldots,n_{r}\},\;\mathbf{j}\in\{1,\ldots,n_{a}\}^{\,r-1}\}$, with radial boundaries $r_{i}=r_{\min}+(r_{\max}-r_{\min})\tfrac{i}{n_{r}},\;i=0,\ldots,n_{r}$, and angular boundaries $\theta_{j,k}=\theta_{\min,k}+(\theta_{\max,k}-\theta_{\min,k})\tfrac{j}{n_{a}},\;j=0,\ldots,n_{a},\;k=1,\ldots,r-1$. The dominant cost of index construction comes from full Mahalanobis-based assignments during optimization, requiring $\mathcal{O}(|X|\cdot K\cdot d^{2})$ time per iteration, while initialization, refinement, and PCA quantization add only lower-order terms (see Appendix A.3).
@@ 2.5 Inference
Given a query $\mathbf{q}$, we first select the top-$k_{G}$ Gaussians by Mahalanobis score $\delta_{M}(\mathbf{q},g)$. For each selected Gaussian $g$, we project $\mathbf{q}$ to its local PCA space $\mathbf{q}^{r}_{g}=\mathbf{P}_{g}^{\top}(\mathbf{q}-\bar{\mathbf{x}}_{g})$ and convert to hyperspherical coordinates $\mathbf{s}_{g}=\text{cart2sph}(\mathbf{q}^{r}_{g})$. We prioritize bins by a query-to-bin distance computed in the reduced Euclidean space. Each bin $B_{i,\mathbf{j}}$ is defined by spherical bounds $r\in[r_{i},r_{i+1}]$ and $\theta_{k}\in[\theta_{j_{k},k},\theta_{j_{k}+1,k}]$ for $k=1,\dots,r-1$. We define:
[…]
If $\mathbf{s}_{g}$ lies inside $B_{i,\mathbf{j}}$, then $d(\mathbf{q}^{r}_{g},B_{i,\mathbf{j}})=0$; otherwise we solve the bound-constrained problem (Byrd et al., 1995) initialized at $\phi_{1}=\text{clip}(\|\mathbf{q}^{r}_{g}\|_{2},[r_{i},r_{i+1}])$ and $\phi_{k+1}=0.5(\theta_{j_{k},k}+\theta_{j_{k}+1,k})$. Bins in $\mathcal{Q}_{g}$ are sorted by $d(\mathbf{q}^{r}_{g},\cdot)$ ascending and scanned until a bin budget $\rho\in(0,1]$ of bins per cell is exhausted (typically $\rho=0.3$). Within each visited bin we compute exact Euclidean distances in the original space between $\mathbf{q}$ and all points indexed in that bin, aggregating candidates across the $k_{G}$ selected Gaussians. The final result is obtained by selecting the overall nearest neighbors among the accumulated candidates, with the dominant inference cost coming from distance computations (see Appendix A.3).
@@
BIRCHGMMOurs 5%Ours 2.5%Ours 1%
(c) Efficiency (Sift)
@@
(e) MNIST
(f) Fashion-MNIST